Walt made an extra $10,000 last year from a part-time job. He invested part of the money at 7% and the rest at 10%. He made a total of $820 in interest. How much was invested at 10%?
Answer by mananth(16949)

Part I 7.00% per annum ------------- Amount invested =x
Part II 10.00% per annum ------------ Amount invested = y
10000
Interest----- 820.00

Part I 7.00% per annum ---x *
Part II 10.00% per annum ---y
Total investment
x + 1 y= 10000 -------------1
Interest on both investments
7.00% x + 10.00% y= 820
Multiply by 100
7 x + 10 y= 82000.00 --------2
Multiply (1) by -7
we get
-7 x -7 y= -70000.00
Add this to (2)
0 x 3 y= 12000
divide by 3
y = 4000 - x)
Part I 7.00% $ 6000
Part II 10.00% $ 4000

CHECK
6000 --------- 7.00% ------- 420.00
4000 ------------- 10.00% ------- 400.00
Total -------------------- 820.00

m.ananth@hotmail.ca